[Brahms-l] Shift report 20050521 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat May 21 2005 - 08:06:08 EDT
Shifter leader was rk
Shifters were 


1. After false fire alarm at 2 o'clock beam injected into RICH at 0400 was quickly dumped which was caused by trip in our D1 magnet. Another injection at 0500, taking data since 0600.
2. Pedestals, 2 physics runs.
3. At 0730 I discovered that CC is not working (for 5-10 min I would say). Reseting BB3 mainframe thrice didn't work, now running without it.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat May 21 08:06:20 2005

This archive was generated by hypermail 2.1.8 : Sat May 21 2005 - 08:06:33 EDT